OpenStudy (anonymous):

w squared+3w-4 can you help me on this

OpenStudy (anonymous):

This how you would factor it. f(x)=(W+4)(W-1) The zeroes of the function =-4 and 1

OpenStudy (phi):

Do you want to factor this? If so, look at the third term (the minus four). Find all the factors of 4: 1,4 2,2 Now notice the sign of 4 is minus. This means the difference between one of the pairs has to be the +3 (from the second term 3w) The second pair 2,2 won't work: 2-2 = 0 (not 3) But 1-4= -3. Except you want +3, so try 4-1= 3 Yes, it works. So the pair you want is 4, -1 and the factors are: (x+4)(x-1) Let's check by multiplying it out: x*x + x*-1+4*x+4*-1= x^2 +3x -4 Yes, it worked.